# Debate-to-Detect: Reformulating Misinformation Detection as a Real-World Debate with Large Language Models

### v3

@@ Abstract
The proliferation of misinformation in digital platforms reveals the limitations of traditional detection methods, which mostly rely on static classification and fail to capture the intricate process of real-world fact-checking. Despite advancements in Large Language Models (LLMs) that enhance automated reasoning, their application to misinformation detection remains hindered by issues of logical inconsistency and superficial verification. In response, we introduce Debate-to-Detect (D2D), a novel Multi-Agent Debate (MAD) framework that reformulates misinformation detection as a structured adversarial debate. Inspired by fact-checking workflows, D2D assigns domain-specific profiles to each agent and orchestrates a five-stage debate process, including Opening Statement, Rebuttal, Free Debate, Closing Statement, and Judgment. To transcend traditional binary classification, D2D introduces a multi-dimensional evaluation mechanism that assesses each claim across five distinct dimensions: Factuality, Source Reliability, Reasoning Quality, Clarity, and Ethics. Experiments with GPT-4o on two datasets demonstrate significant improvements over baseline methods, and the case study highlight D2D’s capability to iteratively refine evidence while improving decision transparency, representing a substantial advancement towards interpretable misinformation detection. The code will be released publicly after the official publication.
@@
Chen Han1,2, Wenzhen Zheng2, Xijin Tang1,2 1 School of Advanced Interdisciplinary Sciences, UCAS 2 State Key Laboratory of Mathematical Sciences, AMSS, CAS {hanchen23, zhengwenzhen21}@mails.ucas.ac.cn, xjtang@iss.ac.cn
@@ 1 Introduction
Figure 2: The D2D framework structures misinformation detection as a multi-agent debate, comprising two layers: the Agent Layer and the Orchestrator Layer. The Orchestrator Layer (a) coordinates the debate process through five stages—Opening, Rebuttal, Free Debate, Closing, and Judgement—while maintaining a shared memory. The Agent Layer (b) comprises domain-specific agents, including the Affirmative, Negative, and Judge roles.
To overcome the constraints, researchers have introduced multi-step reasoning and multi-agent strategies, including Chain-of-Thought (CoT) (Wei et al., 2022), Self-Reflection (Madaan et al., 2023; Shinn et al., 2023), and Multi-Agent Debate (MAD) (Du et al., 2024; Liang et al., 2024; Li et al., 2024; Amayuelas et al., 2024). While these methods have shown efficacy in mitigating hallucinations and enhancing response quality, their evaluations are often restricted to controlled settings with limited contextual diversity, failing to capture the complexity of real-world misinformation (Deng et al., 2025). Moreover, existing MAD frameworks lack the structured process of real-world fact-checking, where claims are systematically examined through evidence collection, counterargument analysis, and multi-dimensional evaluation conducted by domain experts (Slonim et al., 2021; Masterman et al., 2024). Current MAD frameworks capture only fragmented components of the process, relying on general agents and neglecting the differentiation of distinct debate stages, resulting in simplified binary judegments.
In this study, we propose Debate-to-Detect (D2D), an extension of MAD that simulates realistic debate with a set of domain-specific LLM agents. Given an input text, D2D first (i) identifies its topical domain, (ii) assigns each agent a concise domain profile, and (iii) conducts a five-stage debate, including opening statement, rebuttal, free debate, closing statement, and judgement. The judging panel assesses the debate along five independent dimensions, culminating in an authenticity score that reflects both the truthfulness of the claim and the quality of the process. By reformulating misinformation detection as an adversarial debate, our framework enhances the interpretability and better aligns with fact-checking practices.
@@
(1) We introduce Debate-to-Detect (D2D), a structured deliberative framework for misinformation detection inspired by real-world fact-checking workflows. D2D assigns domain-specific profiles to agents, engaging them in a five-stage progressive debate. This structured debate enhances logical coherence and facilitates stepwise evidence refinement, reflecting human reasoning patterns. Experiment results demonstrate that D2D not only significantly outperforms baseline methods but also remains robust on recently published news beyond GPT-4o’s pre-training.
@@ 2 Our Framework: Debate to Detect
Figure 2 illustrates the framework of Debate-to-Detect (D2D), a MAD system that formulates misinformation detection as a structured debate. The framework includes two layers: the Agent Layer, responsible for role profile assignment and task allocation to support diverse argumentation; and the Orchestrator Layer, which manages the debate flow and aggregates judgements.
@@ 2.1 Agent Layer
To enable multi-dimensional evaluation and minimize bias, six judge agents are deployed, each evaluating arguments along specific dimensions. Unlike single-agent evaluations, the multi-judge setup enhances robustness and aligns with the ChatEval strategy for diversified assessment (Chan et al., 2024). Role-specific prompts further promote argumentative diversity and address the "Degeneration-of-Thought" (DoT) issue observed in LLM-based debates (Du et al., 2024).
@@ 2.2 Orchestrator Layer
The Orchestrator Layer organizes the debate into five structured stages: Opening Statement, Rebuttal, Free Debate, Closing Statement, and Judgement. Each stage serves a distinct rhetorical purpose, reinforcing clarity and coherence in argumentation. The Affirmative side initiates the debate, and the Free Debate rounds are configurable to match task-specific requirements.
Before each turn, the active agent receives a compressed summary of the Shared Memory, ensuring engagement with core arguments while filtering redundant information. This mechanism maintains contextual relevance across debate stages.
@@ 2.3 Scoring Mechanism
Following the Closing Statement, the Agent Layer will initiate a two-step judgement process: (1) Neutral Synopsis – A judge agent generates a comprehensive summary of the debate; (2) Scoring – Five independent judge agent assess both sides across the following dimensions: Factuality, Source Reliability, Reasoning Quality, Clarity, and Ethics. Each Judge assigns complementary integer scores summing to 7 (e.g., 4:3, 5:2, 6:1), adhering to a strict zero-sum structure. This design guarantees an unambiguous outcome—since the total score across all dimensions is inherently imbalanced, a tie is mathematically impossible. Consequently, each news is definitively classified as REAL or FAKE.
@@ 3.1 Experimental Setup
  • D2D w/o SD (Multi-dimensional Judgement): This variant eliminates the multi-dimensional judgement mechanism, and a single-dimensional judgement is applied, focusing on the factuality of claims.
Model Configuration. All experiments are conducted using GPT-4o as the backbone model. Domain inference is performed with a temperature of 0.0. Agent profile generation and responses in the four debate stages are generated with a temperature of 0.7 to promote diversity. Judgement is conducted at a lower temperature of 0.2 to ensure consistency and reliability in evaluation. Unless otherwise specified, the number of Free Debate rounds is fixed at 1.
@@ 3.2 Results
We measure the performance using four standard metrics: accuracy, precision, recall, and F1-score. Table 2 presents the overall results of D2D, baselines and ablated variants on the two datasets.
The improvement from ZS to CoT and SR demonstrates a clear improvement in misinformation detection, highlighting the benefits of iterative reasoning mechanisms. Specifically, CoT enhances performance over ZS by approximately 6.74% and 5.98% in F1-score on Weibo21 and FakeNewsDataset, respectively. The SR method further refines these results, achieving 76.00% and 73.40% in F1-score on the two datasets, reflecting the effectiveness of self-evaluation and iterative refinement.
Incorporating adversarial interactions through SMAD results in additional gains, with 76.52% and 74.97% F1-score on Weibo21 and FakeNewsDataset, respectively, representing a small improvement over SR, indicating that structured two-agent debate enhances evidence evaluation by introducing conflicting perspectives.
@@ 4.4 Generalization to Latest Published News
5 Related Work
5.1 Misinformation Detection
5.2 Multi-Agent Debates
@@ 6 Conclusion
In this study, we propose Debate-to-Detect (D2D), a structured debate framework for misinformation detection. By assigning domain-specific profiles to agents based on the inferred topic of the claim, D2D orchestrates a five-stage progressive debate that simulates professional fact-checking workflows. To improve interpretability and analytical rigor, the framework adopts a five-dimensional evaluation method, generating authenticity scores with clear justifications. Experiments on Weibo-21 and FakeNewsDataset show that D2D significantly outperforms baseline methods. Further analysis highlights the importance of agent capability during the judgment stage and reveals that the optimal number of debate rounds varies with text length.
@@
Our work aims to improve the interpretability and robustness of misinformation detection through the Debate-to-Detect (D2D) framework. We acknowledge the ethical considerations inherent in using LLMs for misinformation detection, particularly in environments such as political discourse and public health.
One major concern associated with the use of LLMs in misinformation detection is the potential for biased or erroneous inferences, which can arise from data imbalances or hallucinations. Furthermore, D2D’s agent-driven debate, while designed to simulate human-like argumentation, may not fully replicate the nuance and contextual understanding required for real-world fact-checking, especially in culturally sensitive topics or politically charged narratives.
To address these challenges, our research implements several mitigation strategies. First, D2D applies domain-specific profiles to enhance the contextual relevance. Second, each agent is guided by structured prompts to maintain focus on evidence-based reasoning, minimizing the influence of speculative or unfounded claims. Finally, the judgment stage incorporates a multi-dimensional evaluation mechanism to assess the arguments, ensuring a balanced and transparent decision-making process.
Future work will also focus on enhancing bias detection and correction mechanisms, expanding domain coverage to include underrepresented topics, and incorporating external fact-checking databases to support more reliable judgment. Additionally, we plan to explore more robust fairness evaluations to improve the neutrality and reliability of D2D. We believe that our efforts to refine D2D’s design and evaluation process are essential to mitigating the risks associated with LLM-driven misinformation detection, promoting fair and transparent deliberation across diverse information landscapes.
@@ Appendix A Experiments on full datasets
In this appendix, we present the experimental results on the original datasets. Our main experiments are conducted on preprocessed versions of Weibo21 and FakeNewsDataset to ensure data quality and reduce noise. By evaluating the raw datasets, we illustrate how low-quality and ambiguous samples can degrade the model performance. The statistics of original datasets are presents in Table 6.
@@
Figure 6: Examples of low-quality samples in Weibo21.
A.1 Performance on Full Datasets
@@ Appendix B Prompts Archive
As a public health expert with over 15 years of experience, I have dedicated my career to ensuring the dissemination of accurate and evidence-based health information. I hold a Master’s degree in Public Health from a leading university, supplemented by extensive research in health communication and media studies. My work has involved close collaboration with healthcare professionals, policy makers, and journalists to improve public understanding of health topics. This unique blend of expertise in both health and media allows me to effectively advocate for the crucial role factual news plays in shaping public health outcomes.

### v4

@@ Abstract
The proliferation of misinformation in digital platforms reveals the limitations of traditional detection methods, which mostly rely on static classification and fail to capture the intricate process of real-world fact-checking. Despite advancements in Large Language Models (LLMs) that enhance automated reasoning, their application to misinformation detection remains hindered by issues of logical inconsistency and superficial verification. Inspired by the idea that "Truth Becomes Clearer Through Debate", we introduce Debate-to-Detect (D2D), a novel Multi-Agent Debate (MAD) framework that reformulates misinformation detection as a structured adversarial debate. Based on fact-checking workflows, D2D assigns domain-specific profiles to each agent and orchestrates a five-stage debate process, including Opening Statement, Rebuttal, Free Debate, Closing Statement, and Judgment. To transcend traditional binary classification, D2D introduces a multi-dimensional evaluation mechanism that assesses each claim across five distinct dimensions: Factuality, Source Reliability, Reasoning Quality, Clarity, and Ethics. Experiments with GPT-4o on two fakenews datasets demonstrate significant improvements over baseline methods, and the case study highlight D2D’s capability to iteratively refine evidence while improving decision transparency, representing a substantial advancement towards robust and interpretable misinformation detection. The code will be released publicly after the official publication.
@@
Chen Han1,2, Wenzhen Zheng2, Xijin Tang1,2 1 School of Advanced Interdisciplinary Sciences, University of Chinese Academy of Sciences 2 State Key Laboratory of Mathematical Sciences, Academy of Mathematics and Systems Science, Chinese Academy of Sciences {hanchen23, zhengwenzhen21}@mails.ucas.ac.cn, xjtang@iss.ac.cn
@@ 1 Introduction
To overcome the constraints, researchers have introduced multi-step reasoning and multi-agent strategies, including Chain-of-Thought (CoT) (Wei et al., 2022), Self-Reflection (Madaan et al., 2023; Shinn et al., 2023), and Multi-Agent Debate (MAD) (Du et al., 2024; Liang et al., 2024; Li et al., 2024; Amayuelas et al., 2024). While these methods have shown efficacy in mitigating hallucinations and enhancing reasoning ability, their evaluations are often restricted to controlled settings with limited contextual diversity, failing to capture the complexity of real-world misinformation (Deng et al., 2025). Moreover, existing MAD frameworks lack the structured process of fact-checking, where claims are systematically examined through evidence collection, counterargument analysis, and multi-dimensional evaluation conducted by domain experts (Slonim et al., 2021; Masterman et al., 2024). Current MAD frameworks focus on fragmented elements, employ generic agents, and neglect distinct debate stages, resulting in simplified binary judegments.
Inspired by the idea that “truth becomes clearer through debate,” we propose Debate-to-Detect (D2D), a novel MAD framework that simulates the fact-checking process through structured adversarial debates with LLM agents. Given an input text, D2D (i) identifies its topical domain, (ii) assigns each agent a concise domain profile, and (iii) orchestrates a five-stage debate comprising opening statements, rebuttals, free debate, closing statements, and judgment. A judging panel then evaluates the debate across five independent dimensions, producing an authenticity score that reflects both the truthfulness of the claim and the quality of the reasoning process. By reformulating misinformation detection as debate, D2D achieves higher accuracy while enhancing interpretability, aligning with real-world fact-checking practices.
@@
(1) We introduce D2D, a structured deliberative framework for misinformation detection inspired by real-world fact-checking workflows. D2D assigns domain-specific profiles to agents, engaging them in a five-stage progressive debate. This structured debate enhances logical coherence and facilitates stepwise evidence refinement, reflecting human reasoning patterns. Experiment results demonstrate that D2D not only significantly outperforms baseline methods but also remains robust on recently published news beyond GPT-4o’s pre-training.
@@
Figure 2: The D2D framework structures misinformation detection as a multi-agent debate, comprising two layers: the Agent Layer and the Orchestrator Layer. The Orchestrator Layer (a) coordinates the debate process through five stages—Opening, Rebuttal, Free Debate, Closing, and Judgement—while maintaining a shared memory. The Agent Layer (b) comprises domain-specific agents, including the Affirmative, Negative, and Judge roles. The Judge Agents evaluate the debate along five independent dimensions, and generates both a binary authenticity judgment and a debate summary.
2 Related Work
2.1 Misinformation Detection
2.2 Multi-Agent Debates
@@ 3 Our Framework: Debate to Detect
Figure 2 illustrates the framework of D2D, consisting of two layers: the Agent Layer, which assigns role profiles and allocates tasks to enable diverse argumentation; and the Orchestrator Layer, which controls the debate flow and integrates judgments.
@@ 3.1 Agent Layer
To enable multi-dimensional evaluation and minimize bias, six judge agents are deployed, each evaluating arguments along specific dimensions. Unlike single-agent evaluations, the multi-judge setup enhances robustness and aligns with the ChatEval strategy for diversified assessment (Chan et al., 2024). Role-specific profiles further promote argumentative diversity and relieve the "Degeneration-of-Thought" (DoT) issue observed in LLM-based debates (Du et al., 2024).
@@ 3.2 Orchestrator Layer
The Orchestrator Layer structures the debate into five progressive stages: Opening Statement, Rebuttal, Free Debate, Closing Statement, and Judgement. This progression ensures both breadth and depth in reasoning. The Opening Statement introduces the core arguments from both sides, followed by the Rebuttal, where opposing claims are directly challenged. The Free Debate stage enables agents to flexibly extend, refine, and contest arguments, allowing the debate to evolve beyond scripted exchanges and surface novel reasoning paths. The process concludes with the Closing Statement, which consolidates arguments, and the Judgement, where the multi-judge panel delivers an evidence-based decision.
A key mechanism supporting this structure is the Shared Memory, which accumulates all prior arguments and evidence. Before each turn, the active agent receives a compressed summary of this memory. The summarization constrains generation by highlighting salient arguments, suppressing redundancy, and preserving coherence across stages. This design prevents drift and ensures that agents consistently engage with the central points of contention rather than digressing into irrelevant or repetitive content.
@@ 3.3 Scoring Mechanism
Following the Closing Statement, the Agent Layer will initiate a two-step judgement process: (1) Neutral Synopsis: A judge agent generates a comprehensive summary of the debate; (2) Scoring: Five independent judge agent assess both sides across the following dimensions (Soprano et al., 2021): Factuality, Source Reliability, Reasoning Quality, Clarity, and Ethics. Each Judge assigns complementary integer scores summing to 7 (e.g., 4:3, 5:2, 6:1), adhering to a strict zero-sum structure. This design guarantees an unambiguous outcome—since the total score across all dimensions is inherently imbalanced, a tie is mathematically impossible. Consequently, each news is definitively classified as REAL or FAKE.
@@ 4.1 Experimental Setup
  • BERT Devlin et al. (2019): A fine-tuned BERT-base model for binary classification.
  • RoBERTa Liu et al. (2019): A fine-tuned RoBERTa-base model with the same setup as BERT, serving as a stronger discriminative baseline.
@@
  • D2D w/o MJ (Multi-dimensional Judgement): This variant eliminates the multi-dimensional judgement mechanism, and a single-dimensional judgement is applied, focusing on the factuality of claims.
Model Configuration. All experiments use GPT-4o as the base model. LLM-agents are initialized with predefined prompts provided in Appendix B. Agent response lengths are capped at 1024 tokens. Domain inference and final judgment are conducted with a temperature of 0.0 to ensure stability. To encourage diversity, profile generation and debate responses across all stages use a temperature of 0.7. Unless otherwise specified, the number of Free Debate rounds is fixed at 1.
@@ 4.2 Results
We measure the performance using four standard metrics: accuracy, precision, recall, and F1-score. Table 2 presents the overall results of D2D, baselines and ablated variants on the two datasets. Across all datasets and metrics, D2D achieves the best performance, significantly outperforming both fine-tuned transformers and prompting-based approaches. Although models like RoBERTa achieve competitive performance, they remain limited in interpretability.
The improvement from ZS to CoT and SR demonstrates a clear improvement in misinformation detection, highlighting the benefits of iterative reasoning mechanisms. Specifically, CoT enhances performance over ZS by approximately 6.74% and 5.98% in F1-score on Weibo21 and FakeNewsDataset, respectively. The SR method further refines these results, achieving 76.00% and 73.40% in F1-score on the two datasets, reflecting the effectiveness of self-evaluation and iterative refinement. Incorporating adversarial interactions through SMAD results in additional gains, with 76.52% and 74.97% F1-score on Weibo21 and FakeNewsDataset, respectively, representing a small improvement over SR, indicating that structured two-agent debate enhances evidence evaluation by introducing conflicting perspectives.
@@ 6 Conclusion
In this paper, we introduced Debate-to-Detect (D2D), a structured multi-agent debate framework that reformulates misinformation detection as an adversarial deliberation process. By assigning domain-specific profiles, orchestrating a five-stage debate, and applying a multi-dimensional evaluation rubric, D2D improves both accuracy and interpretability over strong baselines. Our analysis further demonstrates its robustness, generalization beyond memorized content, and resilience to biases such as speaker order and lexical framing. A case study illustrates D2D’s ability to progressively refine evidence and deliver criterion-based evaluations, closely mirroring real-world fact-checking workflows.
Future work will focus on extending D2D to multimodal misinformation (e.g., images, videos, and deepfakes), integrating external fact-checking databases to reduce hallucinations, and enhancing its persuasive capacity—not only classifying claims as true or false, but also explaining why they are misleading. These directions are essential for advancing reliable, transparent, and socially responsible AI systems for misinformation detection.
@@
A major concern in LLM-based misinformation detection is the risk of biased or erroneous inferences, arising from data imbalances or hallucinations. Moreover, D2D’s agent-driven debates, while designed to simulate human argumentation, may fall short in capturing the nuance required for real-world fact-checking, particularly in culturally sensitive or politically charged contexts. These concerns are especially acute in high-stakes domains like law, medicine, and politics, where misleading arguments may erode trust, destabilize communities, or harm individual rights.
@@ Appendix A Experiments on full datasets
In this appendix, we present the experimental results on the original datasets. Our main experiments are conducted on preprocessed versions of Weibo21 and FakeNewsDataset to ensure data quality and reduce noise. By evaluating the raw datasets, we illustrate how low-quality and ambiguous samples can degrade the model performance. The statistics of original datasets are presents in Table 6. The preprocessed datasets are also available at 4open.science/emnlp_d2d-36E2.
@@ Appendix B Prompts Archive
Health: As an experienced health communications expert, I specialize in analyzing and disseminating medical news and information. With a decade of experience working alongside healthcare professionals, researchers, and policymakers, I have a profound understanding of the complexities and dynamics that shape the health news landscape. My work is grounded in ensuring that health news is conveyed accurately and responsibly, leveraging evidence-based research to inform the public. Passionate about promoting health literacy, I am dedicated to enhancing the quality and reliability of health-related news.
Finance: As a financial analyst with extensive experience in equity markets, I focus on evaluating corporate disclosures, earnings reports, and market signals. My expertise lies in assessing financial credibility and detecting inconsistencies across financial statements and media reports. Having collaborated with regulatory agencies and institutional investors, I bring a critical perspective to debates on corporate transparency. I am committed to ensuring that financial information is communicated with clarity and integrity.
Environment: As a climate policy researcher, I have worked with international organizations to assess the impact of environmental regulations on energy sectors. My expertise includes analyzing emission reduction policies, carbon trading mechanisms, and climate adaptation strategies. I bring an evidence-driven approach to discussions of environmental claims, ensuring alignment with the latest scientific findings and policy frameworks. My goal is to promote informed decision-making and constructive dialogue on sustainability issues.